OpenStudy (anonymous):

8|x-3|=88 please tell me step by step

OpenStudy (anonymous):

use the distributive property so 8x(x), and 8x-3 then you have 8x-24=88 move the 4 to the other side8x=88+24 then 8x=112 divide both sides by 8 x=14